The volume of a sphere is increased 72.8%. What is the percentage increase in the radius?
Answer by Boreal(15235)

V is proportional to (4/3) pi*r^3,
Let (4/3) pi=C
V=C*r^3
1.728 V=C (r*^3)
so r*^3/r^3=1.728
take cube root of everything
r*/r=1.2
The percentage increase in the radius is 20%.